OpenStudy (anonymous):

find f''(t) of f(t)= (e^2t i +4tk) X (4t^2j+e^-2t k) X stands for cross product

OpenStudy (amistre64):

product rule is the same for dot or cross

OpenStudy (amistre64):

just make sure you cross the terms and not dot them

OpenStudy (amistre64):

f(t)= (e^2t i +4tk) X (4t^2j+e^-2t k) f'(t)= (e^2t i +4t k)' x (4t^2j+e^-2t k) + (e^2t i +4t k) x (4t^2j+e^-2t k)' f'(t)= (2e^2t i +4 k) x (4t^2j+e^-2t k) + (e^2t i +4t k) x (8t j -2e^-2t k)

OpenStudy (amistre64):

f''(t)= (2e^2t i +4 k)' x (4t^2j+e^-2t k) +(2e^2t i +4 k) x (4t^2j+e^-2t k)' + (e^2t i +4t k)' x (8t j -2e^-2t k) + (e^2t i +4t k) x (8t j -2e^-2t k)' f''(t)= (4e^2t i) x (4t^2j+e^-2t k) +(2e^2t i +4 k) x (8t^ j -2e^-2t k) + (2e^2t i +4 k) x (8t j -2e^-2t k) + (e^2t i +4t k) x (8 j +4e^-2t k)

OpenStudy (anonymous):

thank you :)

OpenStudy (amistre64):

youre welcome :)

OpenStudy (anonymous):

okay so i have to cross all of these and then combine the similar terms right?